OpenStudy (anonymous):

Simplify help please! 1) 5√10 * 3√2 2) 4√3x * 5√12y

OpenStudy (mathmate):

you can use the relations \[\sqrt{a} \sqrt{b} = \sqrt{ab}\] and \[\sqrt{a^{2}bc} = a\sqrt{bc}\]

OpenStudy (anonymous):

I'm just looking for the answers.